# Gaudin models and moduli space of flower curves

### v1

@@ 3.6. Commutative subalgebras of the inhomogeneous Drinfeld-Kohno Lie algebra
$[t_{ij},t_{kl}]=0\text{ for $i,j,k,l$ distinct }\quad[t_{ij},t_{ik}]=-[t_{ij},t_{jk}]\text{ for $i,j,k$ distinct }$
$[t_{ij},u_{k}]=0\text{ for $i,j,k$ distinct }\quad[t_{ij},u_{i}]=-[t_{ij},u_{j}]\quad[u_{i},u_{j}]=0\text{ for $i,j$ distinct }$
@@
Let $Q({\underline{z}})$ be the span of these vectors. This subspace depends only on the image of ${\underline{z}}$ in $F_{n}$ and thus we have a map $F_{n}\rightarrow G(n,{\mathfrak{r}}_{n}^{1})$ defined by ${\underline{z}}\mapsto Q({\underline{z}})$. Note that $Q({\underline{z}})$ always contains $\sum_{i}u_{i}=\sum_{i}h_{i}^{0}({\underline{z}})$ which is central in ${\mathfrak{r}}_{n}$.
@@
We must check that $[h_{i}^{0}({\underline{z}}),h_{j}^{0}({\underline{z}})]=0$. In particular, for every basis element of ${\mathfrak{s}}_{n}^{2}$ (which is $\binom{n}{3}+\binom{n}{2}$ dimensional), we check that the corresponding coefficient is 0. For example, we can consider the coefficient of $[t_{ij},u_{i}]=-[t_{ij},u_{j}]$ which is $\frac{1}{z_{i}-z_{j}}+\frac{1}{z_{j}-z_{i}}=0$. The remaining coefficients are checked in a similar manner. ∎
Let $G_{n}\subset G(n,{\mathfrak{r}}_{n}^{1})$ be the image of $F_{n}$ and let $\overline{G}_{n}$ be its closure in $G(n,{\mathfrak{r}}_{n}^{1})$. Note that every $V\in\overline{G}_{n}$ is commutative since being commutative is a closed condition.
@@
Next, we know that $p_{ij}(Q)\subset{\mathbb{C}}^{3}$ is a commutative subspace (where we identify ${\mathbb{C}}^{3}={\mathfrak{r}}_{2}^{1}$, so $[e_{1},e_{3}]=-[e_{2},e_{3}],\ [e_{1},e_{2}]=0$), which implies that $\dim p_{ij}(Q)\leq 2$.
Finally, we will now find another vector in $p_{ij}(Q)$ by an explicit computation. Since $Q=Q(C)$ for some $C\in\overline{F}_{n}$, we need to find a vector $v\in Q(C)$ such that $p_{ij}(v)$ does not lie in the span of $u_{i}+u_{j}$. Suppose that $C\in W_{\tau}$ for some planar binary tree $\tau$. We have two cases depending on whether $i,j$ lie in the same tree in $\tau$.
If they are in the same tree, then let $v$ be their meet. Assume without loss of generality that $i$ lies in the left branch above $v$. Then $p_{ij}(h_{v}^{0}(C))=(\delta_{pq},0,\delta_{pq}\nu_{ij})$ where $p,q$ are the consecutive leaves meeting at $v$. On $W_{\tau}$, $\delta_{pq}\nu_{ij}$ is non-zero (by [IKLPR, Lemma 6.16]) and so $p_{ij}(h_{v}^{0}(C))$ does not lie in the span of $(1,1,0)$, but does lies in $K(\nu_{ij})$.
On the other hand, if $i,j$ do not lie in the same tree, then let $v$ be the root of the tree containing $i$. Then $p_{ij}(h_{v}^{0}(C))=(1,0,\nu_{ij})$ which again does not lie in the span of $(1,1,0)$, but does lies in $K(\nu_{ij})$. ∎
@@
As in the proof of Lemma 6.17 from [IKLPR], we can find a planar binary forest $\tau$ such that $C\in W_{\tau}$ and the partition of $[n]$ corresponding to the leaves of $\tau$ is exactly ${\mathcal{S}}$. So $\tau$ has trees $\tau_{1},\dots,\tau_{m}$ such that the leaves of $\tau_{r}$ are labelled $S_{r}$. In particular, all $i,j,k$ lies on the same tree.
@@
Since $\mu_{ijk}\nu_{ik}=\nu_{ij}$, we see that $w_{1},w_{2}\in J(\mu_{ijk}(C))$. On the other hand, $\delta_{p_{1}q_{1}}\nu_{ij}\neq 0$ and $\delta_{p_{2}q_{2}}\nu_{ik}\neq 0$ by [IKLPR, Lemma 6.16], so $w_{1},w_{2}$ are linearly independent.
@@ 3.7. Degeneration
$[t_{ij},t_{kl}]=0\text{ for $i,j,k,l$ distinct }\quad[t_{ij},t_{ik}]=-[t_{ij},t_{jk}]\text{ for $i,j,k$ distinct }$
$[t_{ij},u_{k}]=0\text{ for $i,j,k$ distinct }\quad[u_{i},t_{ij}]=-[u_{j},t_{ij}]\quad[u_{i},u_{j}]=-\hbar[u_{i},t_{ij}]\text{ for $i,j$ distinct }$
@@
[…]
@@
Now as before, we let ${\mathcal{G}}_{n}\subset{\mathrm{Gr}}_{{\mathbb{A}}^{1}}(n,{\widetilde{\mathfrak{r}}}^{1}_{n})$ be the image of ${\mathcal{F}}_{n}$ and we consider the closure $\overline{\mathcal{G}}_{n}$. We have $\overline{\mathcal{G}}_{n}\rightarrow{\mathbb{A}}^{1}$ with fibres $\overline{G}_{n}$ over 00 and $\overline{N}_{n+1}$ over $\varepsilon\neq 0$.
We will now define ¯$\overline{\mathcal{F}}_{n}\rightarrow\overline{\mathcal{G}}_{n}$. For every binary planar forest $\tau$ and any internal vertex $v$ of $\tau$, we define $h_{v}^{\hbar}\in\mathcal{O}({\mathcal{W}}_{\tau})\otimes_{\mathbb{C}[\hbar]}{\widetilde{\mathfrak{r}}}^{1}_{n}$ by the formulas (9) and (10). By the same argument as in Lemma 3.9, these generate a rank $n$ subvector bundle of $\mathcal{O}({\mathcal{W}}_{\tau})\otimes_{\mathbb{C}[\hbar]}{\widetilde{\mathfrak{r}}}^{1}_{n}$ and thus we get a map ${\mathcal{W}}_{\tau}\rightarrow{\mathrm{Gr}}_{{\mathbb{A}}^{1}}(n,{\widetilde{\mathfrak{r}}}^{1}_{n})$. By the same argument as in the proof of Proposition 3.10, we get a surjective map ¯$\overline{\mathcal{F}}_{n}\rightarrow\overline{\mathcal{G}}_{n}$.
@@
The morphism ¯$\overline{\mathcal{F}}_{n}\rightarrow\overline{\mathcal{G}}_{n}$ is an isomorphism.
@@
For any $\varepsilon\neq 0$, we can restrict ¯$\overline{\mathcal{F}}_{n}\rightarrow\overline{\mathcal{G}}_{n}$ to a map $\overline{M}_{n+2}\cong\overline{\mathcal{F}}_{n}(\varepsilon)\rightarrow{\mathcal{G}}_{n}(\varepsilon)\cong\overline{N}_{n+1}$. This must agree with the Aguirre-Felder-Veselov isomorphism (Theorem 3.4) since both maps agree on $M_{n+2}$. Combined with Theorem 3.12, we see that for every $\varepsilon\in\mathbb{C}$, the map ¯$\overline{\mathcal{F}}_{n}(\varepsilon)\rightarrow\overline{\mathcal{G}}_{n}(\varepsilon)$ is an isomorphism and hence the whole map is an isomorphism. (This could also be proven without using Theorem 3.4 by an argument similar to the proof of Theorem 3.12.) ∎
@@
which can be extended to derivations $\partial_{t},t\partial_{t}$ of $S(t^{-1}\mathfrak{g}[t^{-1}])$ and $U(t^{-1}\mathfrak{g}[t^{-1}])$. The derivation $t\partial_{t}$ integrates to the loop rotation action of $\mathbb{C}^{\times}$ on $U(t^{-1}\mathfrak{g}[t^{-1}])$ where $x[-k]$ has weight $-k$.
@@
Consider the automorphism $\exp(-u\partial_{t})$ of $S(t^{-1}\mathfrak{g}[t^{-1}])[[u]]$; it maps $x[-1]$ to $\sum\limits_{k=0}^{\infty}x[-k-1]u^{k}=\delta(x)$ for $x\in\mathfrak{g}$. Since these elements generate $S(\mathfrak{g})$, we conclude that $\delta(y)=\exp(-u\partial_{t})i_{-1}(y)$ for any $y\in S(\mathfrak{g})$.
@@ 4.3. Classical homogeneous Gaudin algebra
Note that for $s\in\mathbb{C}^{\times},z\in\mathbb{C}^{\times},a\in U(t^{-1}\mathfrak{g}[t^{-1}])$ we have $\operatorname{ev}_{sz}(a)=\operatorname{ev}_{z}(sa)$ where on the right hand side we use the loop rotation action of $\mathbb{C}^{\times}$ on $U(t^{-1}\mathfrak{g}[t])$. Thus if $a$ has weight $-m$ for this action, then $\operatorname{ev}_{sz}(a)=s^{-m}\operatorname{ev}_{z}(a)$.
@@
At the beginning of this section, we assumed that $z_{i}\neq 0$ for all $i$. This is in fact unnecessary and $A(z_{1},\dots,z_{n})$ is well-defined for all $z_{1},\dots,z_{n}$, pairwise distinct.
@@ 4.4. Definition of the Gaudin subalgebra
From the theorem below we know that there are $12n-$ 1 algebraically independent generators of degree $2$ in $\mathcal{A}(z_{1},\ldots,z_{n})$. Note that $\sum H_{i}=0$ so we can consider $H_{1},\ldots,H_{n-1}$ and $n$ quadratic Casimirs, $\Omega^{(1)},\dots,\Omega^{(n)}$ as a complete list of these generators.
@@
    The subalgebra $\mathcal{A}(z_{1},\ldots,z_{n})$ is polynomial ring with $(n-1)(p+r)+r$ generators. One possible choice of free generators is
@@
Figure 2. This is a curve $C\in\overline{M}_{7+1}$. We have $n=7$, $m=4$, and =($\mathcal{B}=(\{1\},\{2,3\},\{4\},\{5,6,7\})$. Each component contributes the algebra shown and the whole curve gives $\mathcal{A}(C)=\Delta^{\mathcal{B}}(\mathcal{A}(w_{1},w_{2},w_{3},w_{4}))\mathcal{A}(z_{2},z_{3})\mathcal{A}(z_{5},z_{6},z_{7})$.
@@ 5.1. Definition of trigonometric Gaudin subalgebras
We will now a precise description of $\psi$. We begin with the following technical result.
@@
Let $u\in U\mathfrak{n}_{+}$. Then $u\otimes 1-1\otimes\Delta^{n}(S(u))\in U\mathfrak{g}^{\otimes n+1}\Delta^{n+1}(\mathfrak{n}_{+})$ where $S:U\mathfrak{n}_{+}\rightarrow U\mathfrak{n}_{+}$ is the antipode map (the algebra anti-automorphism satisfying $S(x)=-x$ for $x\in\mathfrak{n}_{+})$.
@@
$=(x\otimes 1)(u\otimes 1-1\otimes S(c))+(1\otimes S(c))\Delta(x).$

### v2

@@ 3.6. Commutative subalgebras of the inhomogeneous Drinfeld-Kohno Lie algebra
[…]
@@
Let $Q({\underline{z}})$ be the span of these vectors. This subspace depends only on the image of $\underline{z}$} in $F_{n}$ and thus we have a map $F_{n}\rightarrow{\mathrm{Gr}}(n,{\mathfrak{r}}_{n}^{1})$ defined by ${\underline{z}}\mapsto Q({\underline{z}})$. Note that $Q({\underline{z}})$ always contains $\sum_{i}u_{i}=\sum_{i}h_{i}^{0}({\underline{z}})$ which is central in ${\mathfrak{r}}_{n}$.
@@
We must check that $h_{i}^{0}({\underline{z}}),h_{j}^{0}({\underline{z}})]=0$. In particular, for every basis element of ${\mathfrak{s}}_{n}^{2}$ (which is $\binom{n}{3}+\binom{n}{2}$ dimensional), we check that the corresponding coefficient is 0. For example, we can consider the coefficient of $t_{ij},u_{i}]=-[t_{ij},u_{j}]$ which is $\frac{1}{z_{i}-z_{j}}+\frac{1}{z_{j}-z_{i}}=0$. The remaining coefficients are checked in a similar manner. ∎
Let $G_{n}\subset{\mathrm{Gr}}(n,{\mathfrak{r}}_{n}^{1})$ be the image of $F_{n}$ and let $\overline{G}_{n}$ be its closure in ${\mathrm{Gr}}(n,{\mathfrak{r}}_{n}^{1})$. Note that every $V\in\overline{G}_{n}$ is commutative since being commutative is a closed condition.
@@
Next, we know that $p_{ij}(Q)\subset{\mathbb{C}}^{3}$ is a commutative subspace (where we identify ${\mathbb{C}}^{3}={\mathfrak{r}}_{2}^{1}$, so $e_{1},e_{3}]=-[e_{2},e_{3}],\ [e_{1},e_{2}]=0$), which implies that $\dim p_{ij}(Q)\leq 2$.
Finally, we will now find another vector in $p_{ij}(Q)$ by an explicit computation. Since $Q=Q(C)$ for some $C\in\overline{F}_{n}$, we need to find a vector $v\in Q(C)$ such that $p_{ij}(v)$ does not lie in the span of $u_{i}+u_{j}$. Suppose that $C\in W_{\tau}$ for some planar binary forest $\tau$. We have two cases depending on whether $i,j$ lie in the same tree in $\tau$.
If they are in the same tree, then let $v$ be their meet. Assume without loss of generality that $i$ lies in the left branch above $v$. Then using (10), we have $p_{ij}(h_{v}^{0}(C))=(\delta_{pq},0,\delta_{pq}\nu_{ij})$ where $p,q$ are the consecutive leaves meeting at $v$. On $W_{\tau}$, $\delta_{pq}\nu_{ij}$ is non-zero (by [IKLPR, Lemma 6.16]) and so $p_{ij}(h_{v}^{0}(C))$ does not lie in the span of $(1,1,0)$, but does lie in $K(\nu_{ij})$.
On the other hand, if $i,j$ do not lie in the same tree, then let $v$ be the root of the tree containing $i$. Then using (11), we have $p_{ij}(h_{v}^{0}(C))=(1,0,\nu_{ij})$ which again does not lie in the span of $(1,1,0)$, but does lie in $K(\nu_{ij})$. ∎
@@
As in the proof of Lemma 6.17 from [IKLPR], we can find a planar binary forest $\tau$ such that $C\in W_{\tau}$ and the partition of $[n]$ corresponding to the leaves of $\tau$ is exactly $\mathcal{S}$}. So $\tau$ has trees $\tau_{1},\dots,\tau_{m}$ such that the leaves of $\tau_{r}$ are labelled $S_{r}$. In particular, all $i,j,k$ lie on the same tree.
@@
Recall that
$J(\mu_{ijk})=\bigl\{(x,y,z):x-\mu_{ijk}y+(\mu_{ijk}-1)z=0\bigr\}$
Since $\mu_{ijk}\nu_{ik}=\nu_{ij}$, we see that $w_{1}\in J(\mu_{ijk})$. Also,
$\mu_{ijk}\nu_{ik}+(\mu_{ijk}-1)\nu_{jk}=-\nu_{ij}-\mu_{jik}\nu_{jk}=-\nu_{ij}-\nu_{ji}=0$
and so $w_{2}\in J(\mu_{ijk})$. On the other hand, $\delta_{p_{1}q_{1}}\nu_{ij}\neq 0$ and $\delta_{p_{2}q_{2}}\nu_{ik}\neq 0$ by [IKLPR, Lemma 6.16], so $w_{1},w_{2}$ are linearly independent.
@@ 3.7. Degeneration
[…]
For $\varepsilon\neq 0$, we have an isomorphism ${\widetilde{\mathfrak{r}}}_{n}(\varepsilon)\cong{\mathfrak{s}}_{n+1}$ given $t_{ij}\mapsto t_{ij}$ and $u_{i}\mapsto-\varepsilon t_{i\,0}$. On the other hand, there is an obvious isomorphism ${\widetilde{\mathfrak{r}}}_{n}(0)\cong{\mathfrak{r}}_{n}$.
@@
$\begin{gathered}h^{\varepsilon}_{i}({\underline{z}})=u_{i}-\sum_{j\neq i}\frac{1-\varepsilon z_{i}}{z_{j}-z_{i}}t_{ij}=(1-\varepsilon z_{i})\Bigl(\frac{-\varepsilon}{1-\varepsilon z_{i}}t_{i\,0}-\sum_{j\neq i}\frac{1}{z_{j}-z_{i}}t_{ij}\Bigr)\\ =(1-\varepsilon z_{i})\Bigl(\frac{1}{z_{i}-\varepsilon^{-1}}t_{i\,0}+\sum_{j\neq i}\frac{1}{z_{i}-z_{j}}t_{ij}\Bigr)\end{gathered}$
@@
(12) ${\mathcal{F}}_{n}(\varepsilon)\cong M_{n+2}$${\mathcal{F}}_{n}$$F_{n}$$Gr(n,{\mathfrak{s}}^{1}_{n+1})$${\mathrm{Gr}}_{{\mathbb{A}}^{1}}(n,{\widetilde{\mathfrak{r}}}^{1}_{n})$${\mathrm{Gr}}(n,{\mathfrak{r}}^{1}_{n})$$\{\varepsilon\neq 0\}$${\mathbb{A}}^{1}$$\{0\}$}
Now as before, we let ${\mathcal{G}}_{n}\subset{\mathrm{Gr}}_{{\mathbb{A}}^{1}}(n,{\widetilde{\mathfrak{r}}}^{1}_{n})$ be the image of ${\mathcal{F}}_{n}$ and we consider the closure $\overline{{\mathcal{G}}}_{n}$.
We will now define a morphism $\overline{{\mathcal{F}}}_{n}\rightarrow\overline{{\mathcal{G}}}_{n}$ over ${\mathbb{A}}^{1}$. For every binary planar forest $\tau$ and any internal vertex $v$ of $\tau$, we define $h_{v}^{\hbar}\in\mathcal{O}({\mathcal{W}}_{\tau})\otimes_{\mathbb{C}[\hbar]}{\widetilde{\mathfrak{r}}}^{1}_{n}$ as follows
@@
In the first equation, $i\,\text{left}$” means that $i$ is in the left branch above $v$. In the second equation, $i\,\text{tree}$” means that $i$ is in the tree above $v$. Finally $j\,\text{other}$” means that $j$ lies elsewhere in the forest.
The analog of Lemma 3.8 holds with the same proof; note that $\hbar$ appears in the above formulas for $h_{v}^{\hbar}$ since $\nu_{ij}+\nu_{ji}=\hbar$.
By the same argument as in Lemma 3.9, these generate a rank $n$ subvector bundle of $\mathcal{O}({\mathcal{W}}_{\tau})\otimes_{\mathbb{C}[\hbar]}{\widetilde{\mathfrak{r}}}^{1}_{n}$ and thus we get a map ${\mathcal{W}}_{\tau}\rightarrow{\mathrm{Gr}}_{{\mathbb{A}}^{1}}(n,{\widetilde{\mathfrak{r}}}^{1}_{n})$. By the same argument as in the proof of Proposition 3.10, these glue to a surjective morphism $\overline{{\mathcal{F}}}_{n}\rightarrow\overline{{\mathcal{G}}}_{n}$ (since $\overline{{\mathcal{F}}}_{n}$ is proper over ${\mathbb{A}}^{1}$).
@@
The morphism $\overline{{\mathcal{F}}}_{n}\rightarrow\overline{{\mathcal{G}}}_{n}$ is an isomorphism.
@@
We proceed similar to the proof of Theorem 3.12.
First Lemmas 3.13, and 3.14 generalize to the degeneration in a straightforward way.
For Lemma 3.15, we again define $w_{1}=q_{ijk}(h^{\hbar}_{v_{1}}(C)),w_{2}=q_{ijk}(h^{\hbar}_{v_{2}}(C))$ as in the proof, and we find that
$w_{1}=-(\nu_{ji},\nu_{ki},0)\quad w_{2}=-(\hbar,\nu_{ki},\nu_{kj})$
We observe that
$\nu_{ji}-\mu_{ijk}\nu_{ki}=0$
so $w_{1}\in J(\mu_{ijk})$. On the other hand,
$\hbar-\mu_{ijk}\nu_{ki}+(\mu_{ijk}-1)\nu_{kj}=\hbar-\nu_{ji}-\mu_{jik}\nu_{kj}=\hbar-\nu_{ji}-\nu_{ij}=0$
so also $w_{2}\in J(\mu_{ijk})$. Thus Lemma 3.15 generalizes as well.
Hence we can follow the same argument as the proof of Theorem 3.12 and deduce the result.
@@
which can be extended to derivations $\partial_{t},t\partial_{t}$ of $S(t^{-1}\mathfrak{g}[t^{-1}])$ and $U(t^{-1}\mathfrak{g}[t^{-1}])$. The derivation $t\partial_{t}$ integrates to the loop rotation action of $\mathbb{C}^{\times}$ on $U(t^{-1}\mathfrak{g}[t^{-1}])$ where x[k]x[-k] has weight k-k.
@@
Consider the automorphism $\exp(-u\partial_{t})$ of $S(t^{-1}\mathfrak{g}[t^{-1}])[[u]]$; it maps x[1]x[-1] to $\sum\limits_{k=0}^{\infty}x[-k-1]u^{k}=\delta(x)$ for $x\in\mathfrak{g}$. Since these elements generate $S(\mathfrak{g})$, we conclude that $\delta(y)=\exp(-u\partial_{t})i_{-1}(y)$ for any $y\in S(\mathfrak{g})$.
@@ 4.3. Classical homogeneous Gaudin algebra
Note that for $s\in\mathbb{C}^{\times},z\in\mathbb{C}^{\times},a\in U(t^{-1}\mathfrak{g}[t^{-1}])$ we have $\operatorname{ev}_{sz}(a)=\operatorname{ev}_{z}(sa)$ where on the right hand side we use the loop rotation action of $\mathbb{C}^{\times}$ on $U(t^{-1}\mathfrak{g}[t])$. Thus if $a$ has weight m-m for this action, then $\operatorname{ev}_{sz}(a)=s^{-m}\operatorname{ev}_{z}(a)$.
@@
At the beginning of this section, we assumed that $z_{i}\neq 0$ for all $i$. This is in fact unnecessary and $A(z_{1},\dots,z_{n})$ is well-defined for all $z_{1},\dots,z_{n}$, pairwise distinct, see below Proposition 4.13.
@@ 4.4. Definition of the Gaudin subalgebra
From the theorem below we know that there are 2n12n-1 algebraically independent generators of degree $2$ in $\mathcal{A}(z_{1},\ldots,z_{n})$. Note that $\sum H_{i}=0$ so we can consider $H_{1},\ldots,H_{n-1}$ and $n$ quadratic Casimirs, $\Omega^{(1)},\dots,\Omega^{(n)}$ as a complete list of these generators.
@@
    The subalgebra $\mathcal{A}(z_{1},\ldots,z_{n})$ is polynomial ring with (n1)(p+r)+r(n-1)(p+r)+r generators. One possible choice of free generators is
@@
Figure 3. This is a curve $C\in\overline{M}_{7+1}$. We have $n=7$, $m=4$, and $\mathcal{B}=(\{1\},\{2,3\},\{4\},\{5,6,7\})$. Each component contributes the algebra shown and the whole curve gives $\mathcal{A}(C)=\Delta^{\mathcal{B}}(\mathcal{A}(w_{1},w_{2},w_{3},w_{4}))\mathcal{A}(z_{2},z_{3})\mathcal{A}(z_{5},z_{6},z_{7})$.
@@ 5.1. Definition of trigonometric Gaudin subalgebras
We will now give a precise description of $\psi$. We begin with the following technical result.
@@
Let $u\in U\mathfrak{n}_{+}$. Then $u\otimes 1-1\otimes\Delta^{n}(S(u))\in U\mathfrak{g}^{\otimes n+1}\Delta^{n+1}(\mathfrak{n}_{+})$ where $S:U\mathfrak{n}_{+}\rightarrow U\mathfrak{n}_{+}$ is the antipode map (the algebra anti-automorphism satisfying S(x)=xS(x)=-x for $x\in\mathfrak{n}_{+})$.
@@
$=(x\otimes 1)(u\otimes 1-1\otimes S(u))+(1\otimes S(u))\Delta(x).$